# Conformal-in-the-Loop for Learning with Imbalanced Noisy Data

### v1

@@ 4.5 Example Weighting
Fig. 10: Amount Pruned by Training Step and Alpha.
4.6 Example Pruning
@@
Fig. 11: Mean weight by class across training epochs for the CityScapes dataset at worst value for $\alpha$ = 0.09. Each subfigure shows the distribution of mean class weights at different training epochs, illustrating how the algorithm adjusts class weights over time.
Fig. 12: Imbalanced CIFAR-10 classification performance at 10% noise level. The blue dotted line represents the baseline using Cross Entropy Loss; the red dotted line represents the baseline using Focal Loss. The method outperforms both baselines at all tested values of $\alpha$. At $\alpha=0.20$, the method prunes $\approx$ 2.5% of examples. Performance declines when $\alpha\geq 0.21$.
The introduction of noise adds more uncertainty to the network which reduces the amount of pruning overall. Again, no pruning is done until $\alpha\geq 0.16$; moreover, the amount of pruning is reduced when compared to the 0% noise experiment. Pruning to the amount of noise (10%) is not achieved until approximately $\alpha=0.24$. Fig. 12 shows that the method performs well until $\alpha\geq 0.21$, at which point it starts to degrade in the same fashion as CityScapes experiment explored in Fig. 3. Still, for all values of $\alpha$ until 0.30, the method outperforms both the cross entropy loss and focal loss baselines. At $\alpha=0.20$, the last before the method begins to degrade, the method prunes approximately 2.5% of the examples from training.
(a) Imbalanced CIFAR-10 at noise level 0%
(b) Imbalanced CIFAR-10 at noise level 10%
(c) CityScapes
Fig. 13: Quantile by Training Step and Alpha.
@@ 4.7 Considerations for Selecting $\alpha$
In the context of the CityScapes dataset, this critical threshold is reached with $\alpha$ values of 0.6 and above (see Fig. 13(c)). Notably, this is the point where the method begins to underperform compared to the baseline (see Fig. 3), indicating that $\alpha$ should be chosen to maintain the quantile below this level when dealing with well-labeled, large datasets.
Low values of $\alpha$ can lead the algorithm to overemphasize hard examples, which may decrease the model’s certainty over time (see Fig. 8). This phenomenon is reflected in the quantile value, which one would expect to increase as the network undergoes training. In Fig. 13(a), for $\alpha$ values between 0.10 and 0.15, the quantile appears to plateau; moreover, for $\alpha$ values below 0.10, the quantile declines towards the end of training. This decline is likely a result of the network overfitting to uncertain examples. The impact of this effect becomes more pronounced with the introduction of noise, necessitating even higher values of $\alpha$ to sustain an increasing quantile (see Fig. 13(b)).

### v2

@@ 4.6 Example Pruning
Figure 9: Amount Pruned by Training Step and Alpha.
(a) Imbalanced CIFAR-10 at noise level 0%
(b) Imbalanced CIFAR-10 at noise level 10%
(c) CityScapes
Figure 10: Quantile by Training Step and Alpha.
The introduction of noise adds more uncertainty to the network which reduces the amount of pruning overall. Again, no pruning is done until $\alpha\geq 0.16$; moreover, the amount of pruning is reduced when compared to the 0% noise experiment. Pruning to the amount of noise (10%) is not achieved until approximately $\alpha=0.24$. Fig. 12 shows that the method performs well until $\alpha\geq 0.21$, at which point it starts to degrade in the same fashion as CityScapes experiment explored in Fig. 3. Still, for all values of $\alpha$ until 0.30, the method outperforms both the cross entropy loss and focal loss baselines. At $\alpha=0.20$, the last before the method begins to degrade, the method prunes approximately 2. 5% of the training examples.
@@
Figure 11: Mean weight by class across training epochs for the CityScapes dataset at worst value for $\alpha$ = 0.09. Each subfigure shows the distribution of mean class weights at different training epochs, illustrating how the algorithm adjusts class weights over time.
@@ 4.7 Considerations for Selecting $\alpha$
Figure 12: Imbalanced CIFAR-10 classification performance at 10% noise level. The blue dotted line represents the baseline using Cross Entropy Loss; the red dotted line represents the baseline using Focal Loss. The method outperforms both baselines at all tested values of $\alpha$. At $\alpha=0.20$, the method prunes $\approx$ 2.5% of examples. Performance declines when $\alpha\geq 0.21$.
@@
In the context of the CityScapes dataset, this critical threshold is reached with $\alpha$ values of 0.06 and above (see Fig. 10(c)). Notably, this is the point where the method begins to underperform compared to the baseline (see Fig. 3), indicating that $\alpha$ should be chosen to maintain the quantile below this level when dealing with well-labeled, large datasets.
Low values of $\alpha$ can lead the algorithm to overemphasize hard examples, which may decrease the model’s certainty over time (see Fig. 6(b)). This phenomenon is reflected in the quantile value, which one would expect to increase as the network undergoes training. In Fig. 10(a), for $\alpha$ values between 0.10 and 0.15, the quantile appears to plateau; moreover, for $\alpha$ values below 0.10, the quantile declines towards the end of training. This decline is likely a result of the network overfitting to uncertain examples. The impact of this effect becomes more pronounced with the introduction of noise, necessitating even higher values of $\alpha$ to sustain an increasing quantile (see Fig. 10(b)).